OpenStudy (anonymous):

help??? Line WX contains (−1, 2) and (4, 12) Line YZ contains points (5, 8) and (−2, −6). Lines WX and YZ are perpendicular because the slopes are the same parallel because the product of the slopes is −1 perpendicular because the product of the slopes is −1 parallel because the slopes are the same

OpenStudy (anonymous):

it parallel? but which one e.e

OpenStudy (anonymous):

find the slopes m=(y2-y1)/(x2-x1) if one m = -1/(other m), they are perpendicular if one m = other m they are parallel.

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

Glad it was helpful.